# An Extended Generalized Prandtl-Ishlinskii Hysteresis Model for I2RIS Robot

### v1

@@ 1 Introduction
Unlike conventional rigid robotic manipulators, continuum and snake-like robots present unique challenges in modeling and control (Iqbal et al. (2025)) due to their hysteresis caused by mechanical complexity and nonlinear structural behavior (Esfandiari et al. (2024)). Hysteresis significantly influences accurate modeling and control, making it one of the primary bottlenecks in precision applications. To address this issue, data-driven approaches such as neural networks have been explored. For example, long short-term memory (LSTM) networks have been applied for hysteresis modeling in robotic catheters (Wu et al. (2021)) and continuum robots (Wang et al. (2024)). However, these methods often lack physical meaning and typically require large training datasets, which are not practical for I2RIS. Other classical methods, such as the Bouc-Wen model (Zhang et al. (2023)), the GPI model (Al Janaideh et al. (2010)), and the Modified GPI model (Gao et al. (2023)), have limited flexibility in capturing multi-stage hysteresis behavior.
In particular, our experimental data from the I2RIS robot (see Fig. 3(a)) shows the existence of a multi-stage hysteresis behavior, primarily caused by the cable-driven actuation mechanism. In addition, the small size of the robot, 0.9 mm in diameter and 3 mm in length, restricts the embedding of the sensing modules, causing the accurate position control of the robot to be more challenging (Esfandiari et al. (2024), Esfandiari et al. (2025)). Hence, having an accurate model of the I2RIS robot kinematics is necessary to accurately control the robot and improve patient safety.
Figure 1: The I2RIS attached to the SHER 3.0 handle.
@@
The effectiveness of the proposed method in modeling the hysteresis characteristics of the I2RIS robot, which outperforms the conventional GPI model, is evaluated using experimental results.
@@ 2.1 I2RIS Design
The I2RIS is a 2-DoF snake-like robot, capable of bending ±$\pm 45\degree$ for pitch and yaw DoFs, and includes a micro-gripper mounted on the distal segment with a diameter of 0.9 mm and length of 3.0 mm (Esfandiari et al. (2024)). The actuation unit comprises two DC servo motors (DCX8M, Maxon Motor Inc.) with an integrated reduction gear (GPX8, gear ratio 64:1), encoders (ENX 8MAG 256 pulses/revolution), and controllers (EPOS2 24/2). The low-level motor position/velocity control and the high-level trajectory generation are implemented in C++ and Python, respectively. Communications are handled in ROS1 Noetic. As shown in Fig. 2, an experimental calibration setup was developed to capture the robot’s configuration.
Figure 2: Experimental setup.
@@ 2.2 Extended Generalized Prandtl-Ishlinskii Model
The GPI model (Al Janaideh et al. (2010)) combines a generalized play operator with a density function to characterize hysteresis behavior. The play operator $G_{r}[v](t)$ employs two envelope functions, $\gamma_{l}$ and $\gamma_{r}$, to represent the ascending and descending branches of the hysteresis loop, respectively. The parameter $r$ denotes the magnitude of the backlash, and each GPI play operator is defined for a specific value of $r$ with the form:
Figure 3: (a) An example of I2RIS data with multi-stage hysteresis $S_{1}$,$S_{2}$ and $S_{3}$: Absolute yaw bending angle as a function of absolute yaw motor encoder value. (b) An example of the GPI operator. (c) Output of the GPI model $z_{1}$. (d) Output of the GPI model $z_{2}$.
@@
where $n$ denotes the number of different play operators, and the weighting factor $p(r_{i})$ commonly referred to as the density function. An example of the GPI model output is shown in Fig. 3(a).
@@
This novel switching mechanism allows the EGPI model to capture more than one hysteresis stage that the single GPI model cannot represent.
3 Results
@@ 3.1 Simulation Verification
For simulation verification, we manually selected model parameters (Al Janaideh et al. (2010)) to qualitatively illustrate the behavior of the EGPI model. An input signal $v(t)$ is used to test the output $z(t)$ of the EGPI model, along with the outputs of two underlying GPI models, $z_{1}(t)$, $z_{2}(t)$, which together form $z(t)$. The input signal $v(t)$ is taking the decaying sinusoids wave (Wu et al. (2021)) of the form:
@@
$r_{i}=\begin{cases}0,&\text{for }i=0\\ r_{1}+(i-1)\frac{r_{n}-r_{1}}{n-1},&\text{for }i=1,...,n\\ \end{cases}$ (14)
@@
The GPI model $z_{1}(t)$ is shown in Fig. 3(c) and $z_{2}(t)$ is shown in Fig. 3(d). Since the simulation focuses on illustrating the qualitative input-output relationships, they do not have physical units in Fig. 3(c), (d) and Fig. 4. It can be seen that in the increasing and decreasing parts of the inputs of these two models, only one hysteresis dead zone appears. The final output of the EGPI model $z(t)$ that contains the switching mechanism is shown in Fig. 4. In the increasing part of the input, there are two stages of hysteresis dead zones $S_{1}$ and $S_{2}$; and the same as decreasing part of the input, two stages of hysteresis dead zones are shown in $S_{3}$ and $S_{4}$. The hysteresis stage transition points are characterized by two flag points $v_{f,1}$ and $v_{f,2}$.
@@
Figure 6: Hysteresis modeling results. (a) Yaw motion (input$>$0): yaw angle vs. input. (b) Yaw motion (input$<$0): yaw angle vs. input. (c) Pitch motion (input$>$0): pitch angle vs. input. (d) Pitch motion (input$<$0): pitch angle vs. input. (e) Yaw motion (input$>$0): yaw angle vs. time. (f) Yaw motion (input$<$0): yaw angle vs. time. (g) Pitch motion (input$>$0): pitch angle vs. time. (h) Pitch motion (input$<$0): pitch angle vs. time.
3.2 Hysteresis Modeling and Experimental Results
To capture the hysteresis behavior of the I2RIS in two orthogonal motion planes, we collected data for yaw and pitch motions in a manner similar to that described in our previous work (Esfandiari et al. (2024)). The input encoder value with two directions to the motors are shown as Fig. 5. In the experimental setup (see Fig. 2), the angular configuration of the distal link—specifically the pitch and yaw bending angles is measured by tracking two ArUco markers attached to the sides of a 3D-printed cube mounted on the I2RIS gripper using two microscopes. A total of four datasets are collected, each corresponding to a different motion direction (positive/negative yaw and pitch).
The motion of the I2RIS during increasing and decreasing input encoder values is shown in Fig. 3(a) Due to material hysteresis during tendon elongation, friction between tendons, friction between tendons and robotic joints, as well as other nonlinear effects, it is clearly identified that two distinct dead zones $S_{2}$ and $S_{3}$ appear during the decreasing part of the input (see Fig. 3(a)). To accurately capture the observed hysteresis behavior, the proposed EGPI model is employed. The core idea behind the EGPI modeling approach is to construct an analytical representation of complex, multi-stage hysteresis characteristics. First, because there are two hysteresis stages when the input is decreasing, we design two basic GPI operators $F_{r,1}[v](t)$ and $F_{r,2}[v](t)$ expressed as:
@@ 3.2 Hysteresis Modeling and Experimental Results
where $\gamma_{l}$,$\gamma_{r,1}$ and $\gamma_{r,2}$ are three envelope functions. The function $\gamma_{l}$ is shared by both $H_{r,1}v$ and $H_{r,2}v$, as there is only one dead zone $S_{1}$ at the start of the motion (see Fig. 3(a)). In contrast, $\gamma_{r,1}$ and $\gamma_{r,2}$ are assigned to the two operators respectively, in order to capture the two distinct stages of hysteresis $S_{2}$ and $S_{3}$ during the decreasing phase. The $\kappa$ is acting as a magnitude regulator (Gao et al. (2023)) in order to adjust the dead zone magnitude in $S_{3}$. By taking the weighted summation of the operators, the two GPI model output are expressed as:
@@
The widely adopted density function (Brokate and Sprekels (2012)) in (15) is parameterized by $\lambda$ and $\sigma$. The backlash values $r_{i}$ are uniformly sampled from the interval $[r_{1},r_{n}]$ as specified in (14). One magnitude regulator $\kappa$ is set to adjust the backlash in stage $S_{3}$ (see Fig. 3(a)). To construct the EGPI model, a total of eleven parameters for each motion direction ($a_{1},...,a_{6},\lambda,\sigma,r_{1},r_{n},\kappa$), are determined using the LM method. The objective function is expressed as:
@@
where $J$ is the number of the data, and $\theta_{j}$ is the real measured angle. The resulting EGPI hysteresis model outputs are presented in Fig. 6, along with a comparison with the GPI model results. To focus on the input and output relationships, we correlate the absolute encoder value with the absolute bending angle for better visualization. For example, in Fig. 6(b), the actual motion corresponds to the yaw motor movement with an input encoder value less than zero, as shown in Fig. 5(b). In this case, we use the absolute value of the encoder value to visualize the relationship with the positive yaw bending angle.
To quantitatively evaluate the models performance, three metrics are defined as follows: Root Mean Square Error (RMSE)
@@
Moreover, the EGPI model consistently yields lower RMSE, NRMSE, and MAE values compared to the GPI model. For example, in the yaw motor motion with positive input, the RMSE for EGPI is $0.36^{\circ}$, significantly lower than the GPI model’s RMSE of $2.22^{\circ}$. This reduction in error is particularly evident in the decreasing regions of the input, as shown in Fig. 6(a), where the EGPI model captures the hysteresis behavior more accurately.
@@ 4 Conclusion and future work
In this work, we proposed an EGPI model to effectively capture the hysteresis behavior of I2RIS. Simulation results demonstrated its capability to model multi-stage hysteresis through a novel switching mechanism. Experiments on I2RIS data further validated the model, showing significantly better fitting accuracy than the GPI model across three error metrics. These results highlight the strong potential of the EGPI model to represent more complex hysteresis behaviors in other robotic systems. In future work, we plan to enhance the EGPI model by incorporating the coupling effect between the yaw and pitch motions. Next, we will develop an analytical inverse of the EGPI model to establish a mapping from the I2RIS joint space to the actuation space. With the inverse model, a compensation strategy can be explored to enable precise control of I2RIS motion toward the desired joint angles.

### v2

@@ 1 Introduction
Unlike conventional rigid robotic manipulators, continuum and snake-like robots present unique challenges in modeling and control (Iqbal et al. (2025)) due to their hysteresis caused by mechanical complexity and nonlinear structural behavior (Esfandiari et al. (2024)). Hysteresis significantly influences accurate modeling and control, making it one of the primary bottlenecks in precision applications. To address this issue, data-driven approaches such as neural networks and classical modeling techniques have been explored. For example, long short-term memory (LSTM) networks have been applied for hysteresis modeling in robotic catheters (Wu et al. (2021)) and continuum robots (Wang et al. (2024)). However, these methods often lack physical meaning and typically require large training datasets, which are not practical for I2RIS. Other classical data-driven methods, such as the Bouc-Wen model (Zhang et al. (2023)), the GPI model (Al Janaideh et al. (2010)), and the Modified GPI model (Gao et al. (2023)), have limited flexibility in capturing multi-stage hysteresis behavior.
In particular, our experimental data from the I2RIS robot (see Fig. 2(a)) shows the existence of a multi-stage hysteresis behavior, primarily caused by the cable-driven actuation mechanism. In addition, the small size of the robot, restricts the embedding of the sensing modules, causing the accurate position control of the robot to be more challenging (Esfandiari et al. (2024). Hence, having an accurate model of the I2RIS hysteresis is necessary to accurately control the robot and improve patient safety.
Figure 1: Overview of the I2RIS robot. (a) The I2RIS robot and its dexterous distal unit. (b) Conceptual design of dexterous distal unit. (c) Cross-section of the actuation unit. (d) Cross-section of the instrument unit.
@@
The effectiveness of the proposed method in modeling the hysteresis characteristics of the I2RIS robot, is evaluated using experimental results.
@@ 2.1 I2RIS Design
The I2RIS (see Fig. 1(a)) is a 2-DoF snake-like robot, capable of bending $\pm 45\degree$ for pitch and yaw directions, and equipped with a micro-gripper mounted on the distal segment (0.9 mm diameter and 3.0 mm length) (Esfandiari et al. (2024)). Fig. 1(b) shows the conceptual design of the dexterous distal unit of the I2RIS. The dexterous distal unit is composed of 12 disk-like elements, providing 2-DoF bending joints actuated by four wires. The wire driven mechanism is actuated by the rotational motion of pulleys, whose cross-sections are shown in Fig. 1(c) and Fig. 1(d). Further details on the dexterous distal unit and the drive unit design can be found in Jinno and Iordachita (2021).
@@ 2.2 Extended Generalized Prandtl-Ishlinskii Model
The GPI model (Al Janaideh et al. (2010)) combines a generalized play operator with a density function to characterize hysteresis behavior. The play operator $G_{r}[v](t)$ (see Fig. 2(b)) employs two envelope functions, $\gamma_{l}$ and $\gamma_{r}$, to represent the ascending and descending branches of the hysteresis loop, respectively. The parameter $r$ denotes the magnitude of the backlash, and each GPI play operator is defined for a specific value of $r$ with the form:
Figure 2: (a) An example of I2RIS data with multi-stage hysteresis $S_{1}$$S_{2}$ and $S_{3}$: Absolute yaw bending angle as a function of absolute yaw motor encoder value. (b) An example of the GPI operator. (c) Output of the GPI model $z_{1}$. (d) Output of the GPI model $z_{2}$.
@@
where $n$ denotes the number of different play operators, and the weighting factor $p(r_{i})$ commonly referred to as the density function. An example of the GPI model output is shown in Fig. 2(c).
@@
This novel switching mechanism allows the EGPI model to capture more than one hysteresis stage that the single GPI model fails to represent.
@@ 2.3 Simulation Verification
For simulation verification, we manually selected model parameters (Al Janaideh et al. (2010)) to qualitatively illustrate the performance of the EGPI model. An input signal $v(t)$ is used to test the output $z(t)$ of the EGPI model, along with the outputs of two underlying GPI models, $z_{1}(t)$, $z_{2}(t)$, which together form $z(t)$. The input signal $v(t)$ is taking the decaying sinusoids wave (Wu et al. (2021)) of the form:
@@
[…]
@@
The GPI model $z_{1}(t)$ is shown in Fig. 2(c) and $z_{2}(t)$ is shown in Fig. 2(d). Since the simulation focuses on illustrating the qualitative input-output relationships, they do not have physical units in Fig. 2(c), (d) and Fig. 3. It can be seen that in the increasing and decreasing parts of the inputs of these two GPI models, only one hysteresis dead zone appears. The final output of the EGPI model $z(t)$ that contains the switching mechanism is shown in Fig. 3. In the increasing part of the input, there are two stages of hysteresis dead zones $S_{1}$ and $S_{2}$; and the same as decreasing part of the input, two stages of hysteresis dead zones are shown in $S_{3}$ and $S_{4}$. The hysteresis stage transition points are characterized by two flag points $v_{f,1}$ and $v_{f,2}$.
@@
2.4 Hysteresis Modeling Using EGPI Model
The motion of the I2RIS during increasing and decreasing input encoder values is shown in Fig. 2(a). Due to material hysteresis during tendon elongation, friction between tendons, friction between tendons and robotic joints, as well as other nonlinear effects, it is clearly identified that two distinct dead zones $S_{2}$ and $S_{3}$ appear during the decreasing part of the input (see Fig. 2(a)). To accurately capture the observed hysteresis behavior, the proposed EGPI model is employed. The core idea behind the EGPI modeling approach is to construct an analytical representation of complex, multi-stage hysteresis characteristics. First, because there are two hysteresis stages when the input is decreasing, we design two basic GPI operators $F_{r,1}[v](t)$ and $F_{r,2}[v](t)$ expressed as:
@@ 2.4 Hysteresis Modeling Using EGPI Model
where $\gamma_{l}$$\gamma_{r,1}$ and $\gamma_{r,2}$ are three envelope functions. The function $\gamma_{l}$ is shared by both $H_{r,1}v$ and $H_{r,2}v$, as there is only one dead zone $S_{1}$ at the start of the motion (see Fig. 2(a)). In contrast, $\gamma_{r,1}$ and $\gamma_{r,2}$ are assigned to the two operators respectively, in order to capture the two distinct stages of hysteresis $S_{2}$ and $S_{3}$ during the decreasing phase. The $\kappa$ is acting as a magnitude regulator (Gao et al. (2023)) in order to adjust the dead zone magnitude in $S_{3}$. By taking the weighted summation of the operators, the two GPI model output are expressed as:
@@
The widely adopted density function (Brokate and Sprekels (2012)) in (15) is parameterized by $\lambda$ and $\sigma$. The backlash values $r_{i}$ are uniformly sampled from the interval $r_{1},r_{n}]$ as specified in (14). One magnitude regulator $\kappa$ is set to adjust the backlash in stage $S_{3}$ (see Fig. 2(a)). To construct the EGPI model, a total of eleven parameters for each motion direction ($a_{1},...,a_{6},\lambda,\sigma,r_{1},r_{n},\kappa$), are determined using the LM method. The objective function is expressed as:
@@
where $J$ is the number of the data, and $\theta_{j}$ is the real measured angle.
Figure 4: Experimental setup and bending angle definition.
Figure 6: Hysteresis modeling results. (a) Yaw motion (input>>0): yaw angle vs. input. (b) Yaw motion (input<<0): yaw angle vs. input. (c) Pitch motion (input>>0): pitch angle vs. input. (d) Pitch motion (input<<0): pitch angle vs. input. (e) Yaw motion (input>>0): yaw angle vs. time. (f) Yaw motion (input<<0): yaw angle vs. time. (g) Pitch motion (input>>0): pitch angle vs. time. (h) Pitch motion (input<<0): pitch angle vs. time.
3 Experimental Results
To capture the hysteresis behavior of the I2RIS, we developed an experimental setup (see Fig. 4) to collect data for two orthogonal motion planes: yaw and pitch in a manner similar to that described in our previous work (Esfandiari et al. (2024)). The actuation unit comprises two DC servo motors (DCX8M, Maxon Motor Inc.) with an integrated reduction gear (GPX8, gear ratio 64:1), encoders (ENX 8MAG 256 pulses/revolution), and controllers (EPOS2 24/2). The low-level motor position control and the high-level command generation are implemented in C++ and Python, respectively. Communications are handled in ROS1 Noetic. In the experimental setup, the pitch and yaw bending angles of the dexterous distal unit are measured by tracking two ArUco markers attached to the sides of a 3D-printed cube mounted on the gripper using two microscopes.
A total of four datasets were collected, each corresponding to a different motion direction (positive/negative yaw and pitch). The input encoder signals for both directions of a single motor are shown in Fig. 5.
The resulting EGPI hysteresis model outputs are presented in Fig. 6, along with a comparison with the GPI model results. To focus on the input and output relationships, we correlate the absolute encoder value with the absolute bending angle for better visualization. For example, in Fig. 6(b), the actual motion corresponds to the yaw motor movement with an input encoder value less than zero, as shown in Fig. 5(b). In this case, we use the absolute value of the encoder value to visualize the relationship with the positive yaw bending angle. To quantitatively evaluate the models performance, three metrics are defined as follows: Root Mean Square Error (RMSE)
@@ 3 Experimental Results
Moreover, the EGPI model consistently yields lower RMSE, NRMSE, and MAE values compared to the GPI model. For example, in the yaw motor motion with positive input, the RMSE for EGPI is $0.36^{\circ}$, significantly lower than the GPI model’s RMSE of $2.22^{\circ}$. This reduction in error is particularly evident in the decreasing regions of the input, as shown in Fig. 6(a), where the EGPI model captures the hysteresis behavior more accurately. The main reason for this improvement is that the EGPI model can switch its hysteresis representation across different dead-zone stages, a capability that the GPI model lacks.
@@ 4 Conclusion and future work
In this work, we proposed an EGPI model to effectively capture the hysteresis behavior of I2RIS. Simulation results demonstrated its capability to model multi-stage hysteresis through a novel switching mechanism. Experiments on I2RIS data further validated the model, showing significantly better fitting accuracy than the GPI model across three error metrics. These results highlight the strong potential of the EGPI model to represent more complex hysteresis behaviors in other robotic systems. In future work, we plan to enhance the EGPI model by introducing mathematical constraints to ensure continuity across switching points between submodels. Next, we will develop an analytical inverse of the EGPI model to establish a mapping from the I2RIS joint space to the actuation space. With the inverse model, a compensation strategy can be explored to enable precise control of I2RIS motion toward the desired joint angles.